# Fusion of Movement and Naïve Predictions for Point Forecasting in Univariate Random Walks

### v3

@@ Abstract
Traditional methods for point forecasting in univariate random walks often fail to surpass naïve benchmarks due to data unpredictability. This study introduces a novel forecasting method that fuses movement prediction (binary classification) with naïve forecasts for accurate one-step-ahead point forecasting in univariate random walks. The method’s efficacy is demonstrated through theoretical analysis, simulations, and real-world data experiments. It reliably outperforms naïve forecasts with moderate movement prediction accuracies, such as 0.55, and is superior to baseline models such as the ARIMA, linear regression, MLP, and LSTM networks in forecasting the S&P 500 index and Bitcoin prices. This method is particularly advantageous when accurate point predictions are challenging but accurate movement predictions are attainable, translating movement predictions into point forecasts in random walk contexts.
@@ 1 Introduction
This study addresses the challenge of random walk point forecasting from a novel perspective, focusing on translating movement predictions into point forecasts, the potential of which remains less investigated. By introducing a sign representing the movement direction of the future value, the random error term in the definition of random walks, which is independent and identically distributed (i.i.d.), can be expressed as a positive i.i.d. random error multiplied by this sign. With regard to prediction, the sign can be estimated via movement prediction, and the estimation of the positive error for the future point can be obtained empirically based on both the estimation of movement prediction accuracy and the residual of the naïve forecast on the historical data. Therefore, from this variant definition of random walks, the one-step-ahead point forecasting of a univariate random walk can theoretically be derived from the fusion of movement and naïve predictions. More importantly, this fusion method does not require a loss function for optimization. Instead, it can be optimized by estimating movement prediction accuracy on the validation set, avoiding rolling back to a traditional regression method that often faces bottlenecks in random walk forecasting. Furthermore, movement prediction can be achieved using various machine learning and deep learning algorithms, which are advancing rapidly.
@@ 2.2 Movement prediction of random walks
Various machine learning and deep learning models that utilize a range of advanced techniques have been proposed for movement prediction. Classical and machine learning models employ technical indicators, price data, and fundamental data for movement prediction Huynh et al. (2017); Raza (2017); Sezer et al. (2017), whereas neural networks integrated with selective trading strategies have been shown to be effective in movement prediction Das et al. (2018); Yang et al. (2017). Additionally, LSTM and gated recurrent unit (GRU) networks, incorporating technical indicators such as moving average convergence divergence (MACD), demonstrate their utility in handling movement estimation for stock markets Nelson et al. (2017); Troiano et al. (2018).
@@ 3.2 Theoretical analysis of the fusion results
Based on the above analysis, interestingly, the optimal value of $c$ depends only on $ACC^{\text{movement}}$. Given that when $AC{{C}^{movement}}\in(0.5,1)$, the movement prediction is meaningful, we can obtain the corresponding optimal value of $c$, which falls within the range between 0 and 1. If we set a large value for $c$, the threshold of $ACC^{\text{movement}}$ for improvement from the naïve forecast must also be high. Based on (11), when the $ACC^{\text{movement}}$ is only 0.51, setting a small value for $c$, such as 0.01, can theoretically ensure that the fusion forecast outperforms the naïve forecast in terms of the MSE. The value assigned to the weight $c$ for the fusion prediction represents a trade-off between the likelihood of outperforming the naïve forecast and the magnitude of the improvement. When $c$ is fixed, an increase in $ACC^{\text{movement}}$ directly increases the discrepancy in the MSE between the naïve and fusion forecasts. In a real forecasting scenario where $ACC^{\text{movement}}$ is unknown, we can use the movement prediction accuracy on the training set or validation set as an estimation of $ACC^{\text{movement}}$ and set $c$ based on (11). Because the RMSE is the root of the MSE, obtaining the requirement for $RMS{{E}^{fusion}}<RMS{{E}^{naive}}$ is equivalent to obtaining that for $MS{{E}^{fusion}}<MS{{E}^{naive}}$.
@@ 4.1 Experiment 1: Simulated data
Figure 4 illustrates the performance of the fusion forecasts across various accuracy levels using different fixed weights. As shown in the first four plots, under a smaller value of $c$, such as 0.01 or 0.1, the fusion forecast generally outperforms the naïve forecast, albeit at the expense of potential gains from higher $c$. Conversely, setting a larger $c$ could lead to inferior performance to the naïve forecast if the actual $ACC^{\text{movement}}$ is poor.
@@ 5 Discussion
From the decision fusion perspective, integrating diverse predictions is advantageous because each model contributes unique insights and enhances predictive capabilities Bishop and Nasrabadi (2006). Previously, effectively combining predictions from various models that produce different types of data, such as binary and continuous outputs, has presented a significant challenge. The fusion method proposed in this study is an effective approach for converting binary classification into continuous values, thereby enabling the fusion of diverse data types. This conversion facilitates an efficient integration process, allowing for the comprehensive use of available forecasts to improve decision making in real forecasting scenarios.
Notably, the dataset used for experiments may not be large enough to meet the assumptions made in the theoretical analysis. In addition, while the definition of a random walk adopted in this study assumes i.i.d. increments for each time step, real-world data can be drifted random walks and are subject to measurement errors. All these factors could explain the discrepancies between the theoretical and empirical fusion results. Moreover, the random increments are supposed to be the same for all future time steps, which may not always provide a reliable estimation owing to the fluctuating nature of real-world data.

### v4

@@ Abstract
Point forecasting in univariate random walks is an important but challenging research topic that has attracted numerous researchers. Unfortunately, traditional regression methods for this task often fail to surpass naïve benchmarks due to data unpredictability. From a decision fusion perspective, this study proposes a novel forecasting method, which is derived from a variant definition of random walks, where the random error term for the future value is expressed as a positive random error multiplied by a direction sign. This method, based on the fusion of movement and naïve predictions, does not require a loss function for optimization and can be optimized by estimating movement prediction accuracy on the validation set. This characteristic prevents the fusion method from reverting to traditional regression methods and allows it to integrate various machine learning and deep learning models for movement prediction. The method’s efficacy is demonstrated through simulations and real-world data experiments. It reliably outperforms naïve forecasts with moderate movement prediction accuracies, such as 0.55, and is superior to baseline models such as the ARIMA, linear regression, MLP, and LSTM networks in forecasting the S&P 500 index and Bitcoin prices. This method is particularly advantageous when accurate point predictions are challenging but accurate movement predictions are attainable, translating movement predictions into point forecasts in random walk contexts.
@@ 1 Introduction
This study addresses the challenge of random walk point forecasting from a novel perspective, focusing on translating movement predictions into point forecasts, the potential of which remains less investigated. By introducing a sign representing the movement direction of the future value, the random error term in the definition of random walks, which is independent and identically distributed (i.i.d.), can be expressed as a positive i.i.d. random error multiplied by this sign. With regard to prediction, the sign can be estimated via movement prediction, and the estimation of the positive error for the future point can be obtained empirically based on both the estimation of movement prediction accuracy and the residual of the naïve forecast on the historical data. Therefore, from this variant definition of random walks, the one-step-ahead point forecasting of a univariate random walk can theoretically be derived from the fusion of movement and naïve predictions. More importantly, this fusion method does not require a loss function for optimization. Instead, it can be optimized by estimating movement prediction accuracy on the validation set, avoiding reverting to traditional regression methods, which often face bottlenecks in random walk forecasting. Furthermore, movement prediction can be achieved using various machine learning and deep learning algorithms, which are advancing rapidly.
@@ 2.2 Movement prediction of random walks
Various machine learning and deep learning models that utilize a range of advanced techniques have been proposed for movement prediction. Classical and machine learning models employ technical indicators, price data, and fundamental data for movement prediction Raza (2017); Sezer et al. (2017), whereas neural networks integrated with selective trading strategies have been shown to be effective in movement prediction Das et al. (2018); Yang et al. (2017). Additionally, LSTM and gated recurrent unit (GRU) networks, incorporating technical indicators such as moving average convergence divergence (MACD), demonstrate their utility in handling movement estimation for stock markets Nelson et al. (2017); Troiano et al. (2018).
@@ 3.2 Theoretical analysis of the fusion results
Based on the above analysis, interestingly, the optimal value of $c$ depends only on $ACC^{\text{movement}}$. Given that when $AC{{C}^{movement}}\in(0.5,1)$, the movement prediction is meaningful, we can obtain the corresponding optimal value of $c$, which falls within the range between 0 and 1. If we set a large value for $c$, the threshold of $ACC^{\text{movement}}$ for improvement from the naïve forecast must also be high. Based on (11), when the $ACC^{\text{movement}}$ is only 0.51, setting a small value for $c$, such as 0.01, can theoretically ensure that the fusion forecast outperforms the naïve forecast in terms of the MSE. When the $ACC^{\text{movement}}$ is equal to 1, the optimal value of $c$ is 1. Under these conditions, we can achieve the maximum potential improvement from the fusion process, which is given by the value of $\bar{\epsilon}^{2}$. This improvement represents the full potential of the fusion method. The value assigned to the weight $c$ for the fusion prediction represents a trade-off between the likelihood of outperforming the naïve forecast and the magnitude of the improvement. When $c$ is fixed, an increase in $ACC^{\text{movement}}$ directly increases the discrepancy in the MSE between the naïve and fusion forecasts. In a real forecasting scenario where $ACC^{\text{movement}}$ is unknown, we can use the movement prediction accuracy on the training set or validation set as an estimation of $ACC^{\text{movement}}$ and set $c$ based on (11). Because the RMSE is the root of the MSE, obtaining the requirement for $RMS{{E}^{fusion}}<RMS{{E}^{naive}}$ is equivalent to obtaining that for $MS{{E}^{fusion}}<MS{{E}^{naive}}$.
@@ 4.1 Experiment 1: Simulated data
Figure 4 illustrates the performance of the fusion forecasts across various accuracy levels using different fixed weights. As shown in the first four plots, under a smaller value of $c$, such as 0.01 or 0.1, the fusion forecast generally outperforms the naïve forecast, albeit at the expense of potential gains from higher $c$. Conversely, setting a larger $c$ could lead to inferior performance to the naïve forecast if the actual $ACC^{\text{movement}}$ is poor. As long as $ACC^{\text{movement}}$ is better than 0.5, setting a small value to weight $c$ can ensure that the fusion result has an improvement compared to the naïve forecast, regardless of the actual value of $ACC^{\text{movement}}$.
@@ 5 Discussion
From the simulations of the fusion results, it is observed that among the multiple trials of the fusion process at moderate accuracy levels, such as 0.55 and 0.60, some trials still underperformed compared to the naïve forecasts. However, the significant discrepancy between the median results of these trials and the naïve forecasts suggests that adopting an ensemble approach after each fusion process can achieve a more robust improvement over the naïve forecast.
Furthermore, the proposed fusion method can be viewed as a classification-to-regression conversion in the context of univariate time series forecasting. This is particularly useful when integrating diverse predictions is advantageous, but effectively combining predictions from models that produce different types of data, such as binary and continuous outputs, is challenging Bishop and Nasrabadi (2006); Zhang et al. (2022). The fusion method in this study effectively converts binary classifications into continuous values, enabling the integration of diverse data types. This conversion facilitates an efficient integration process, allowing for comprehensive use of available forecasts to improve decision-making in real forecasting scenarios.
Notably, the dataset used for experiments may not be large enough to meet the assumptions made in the theoretical analysis. In addition, while the definition of a random walk adopted in this study assumes i.i.d. increments for each time step, real-world data can be drifted random walks and are subject to measurement errors. All these factors could explain the discrepancies between the theoretical and empirical fusion results. Moreover, this study assumes the random increments to be the same for all future time steps, which may not always provide a reliable estimation owing to the fluctuating nature of real-world data.